# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v2

@@ V-A3 Metrics
Researchers primarily assess the model’s performance based on two key aspects: navigation and question answering. In the navigation aspect, many researchers adhere to the approach introduced by Das et al[242]. and utilize indicators like the distance to the target object upon completion of navigation ($d_{T}$), the change in distance to target from initial to final position ($d_{\varDelta}$) and the smallest distance to the target at any point in the episode ($d_{min}$) to evaluate the performance of the model. They are tested at 10, 30, or 50 actions away from the target. There are also works that measure it based on indicators such as trajectory length, intersection-over-union score for target object ($IoU$), etc. Regarding question answering, the evaluation mainly involves mean rank ($MR$) of the ground-truth answer in the answer list and accuracy of the answers. Recently, Majumdar et al.[248] introduce the concept of an aggregate LLM-based correctness metric (LLM-Match) to evaluate the accuracy of open-vocabulary answers. Additionally, they assess efficiency by incorporating the normalized length of the agent’s path as a weight for the correctness metric.
@@ V-A4 Limitations
1) Dataset: Constructing datasets requires substantial manpower and resources. Additionally, there are still few large-scale datasets, and the metrics for evaluating model performance vary across different datasets, complicating the testing and comparison of performance, 2) Model: Despite the advancements brought by LLMs, the performance of these models still lags significantly behind human levels. Future work may focus more on effectively storing environmental information explored by agents and guiding them to plan actions based on environmental memory and questions, while also enhancing the interpretability of their actions.
@@ V-B Embodied Grasping
V-B1 Gripper
V-B2 Datasets
Recently, a substantial number of grasping datasets [266, 265, 270, 269, 271] have been generated. These datasets typically contain annotated grasping data based on images (RGB, depth), point clouds, or 3D scenes. With the advent of multimodal large models and the application of foundational language models to robotic grasping, there is an urgent need for datasets that include linguistic text. Consequently, existing datasets have been extended or reconstructed to create semantic-grasping datasets [272, 273, 274, 275]. These datasets are instrumental in studying grasping models grounded in language, enabling agents to develop a broad understanding of semantics.
@@ V-B2 Datasets
TABLE X: Embodied grasping datasets.
@@ V-B3 Language-guided grasping
Spatial reasoning [272] refers to instructions that may include the spatial relationship of the object or part to be grasped, necessitating the inference of grasping posture based on the spatial relationships of objects within the scene. For example, “Grasp the keyboard that is to the right of the brown kleenex box” involves understanding and inferring the spatial arrangement of objects.
Logical reasoning [273], on the other hand, involves instructions that may contain logical relationships requiring inference to discern human intent and subsequently grasp the target. For instance, “I am thirsty, can you give me something to drink?” would prompt the agent to potentially hand over a glass of water or a bottle of a beverage. The agent must ensure that the liquid does not spill during the handover, thus generating a reasonable grasping posture.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Initially, Symbolic Agents, rooted in symbolic reasoning, and Reactive Agents, known for their rapid responsiveness, were widely utilized. However, these agents were limited in handling complex strategies under uncertainty. Learning-based agents were subsequently developed to mitigate this limitation, yet they remained inadequate for large-scale real-world problems. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their problem-solving capabilities from virtual space to the physical world, thereby becoming Embodied Agents [280].
To enable embodied agents to operate in the information-rich and complex real world, the Embodied Multimodal Foundation Model has been developed to provide these agents with multimodal perception and reasoning capabilities. To complete a task, embodied agents typically involves the following process: 1) decomposing the abstract and complex task into specific subtasks, which is referred to as high-level Embodied Task Planning. 2) gradually implementing these subtasks by effectively utilizing Embodied Perception and Embodied Interaction models or leveraging the Foundation Model’s policy function, named low-level Embodied Action Planning. It is worth noting that task planning involves thinking before acting, and is therefore typically considered in cyber space. In contrast, action planning must account for effective interaction with the environment and feedback on this information to the task planner to adjust task planning. Thus, it is crucial for embodied agents to align and generalize their abilities from the cyber space to the physical world.
@@ VI-B1 Planning utilizing the Emergent Capabilities of LLMs
To increase the success rate of these solutions, some methods provide more reference examples as context, such as pre-including a successful plan in the prompt [301]. The semantic relevance of these examples also affects task planning accuracy [302]. Therefore, methods like LLM-Planner have been developed to retrieve the most appropriate examples as context using KNN to find task-type similar and semantically relevant examples [303]. Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [304, 305, 306] . Interestingly, improving chain-of-thought (CoT) reasoning [307] is also a viable approach. ReAct [308] incorporates CoT reasoning with plan generation, providing more logical action generation. And some works utilize code as the reasoning medium instead of natural language. The Chain of Code method is particularly suitable for embodied agents in reasoning and executing tasks, where task planning is generated as code based on the available API library[309, 310, 311]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning, a focus of many LLM-based agent studies. For instance, Socratic Models [312] and Socratic Planner [313] use Socratic questioning to derive reliable planning.
@@ VI-B3 Planning utilizing the VLMs
In the EmbodiedGPT framework, the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning [286]. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [325]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [13, 284] and PaLM-E [283] to achieve alignment of visual and textual features in embodied scenarios. Additionally, innovative research such as Matcha [326] explores the integration of large language models (LLMs) with multimodal perception modules (e.g., weight, touch, sound) to enhance robot planning capabilities. For example, when executing the task ”pick up a plastic block,” the robot can assess its weight by lifting it, determine if the sound it makes when tapped matches expected plastic sounds, or use tactile feedback to sense the block’s hardness or softness. Moreover, the VLP [327] model enhances planning by considering historical information and predicting future visual states. Once an LLM generates an action, the video model can simulate multiple potential video representations for each action. These video models then serve as heuristic functions to evaluate the proximity of each action to the goal, thereby improving planning effectiveness.
@@ VI-C2 Action utilizing VLA model
This paradigm leverages the capabilities of embodied multimodal foundation models for planning and executing actions, unlike the previous approach, where task planning and action execution are performed within the same system, reducing communication latency and improving system response speed and efficiency. In VLA models, the tight integration of perception, decision-making, and execution modules allows the system to handle complex tasks and adapt to changes in dynamic environments more efficiently. This integration also facilitates real-time feedback, enabling the agent to self-adjust strategies, thereby enhancing the robustness and adaptability of task execution [3, 286, 285]. However, this paradigm is undoubtedly more complex and costly, particularly when dealing with intricate or long-term tasks. Additionally, a key issue is that an action planner, without an embodied world model, cannot simulate physical laws using only the internal knowledge of an LLM. This limitation hinders the agent to accurately and effectively complete various tasks in the physical world, preventing the seamless transfer from cyber space to physical world.
@@ VII-A Embodied World Model
TABLE XI: Summary of the embodied world methods discussed in VII-A.
VII-A1 Generation-based Methods
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim2real approach[347], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen[341], Liu et al.[342], Lionel et al. [340]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with large language models (LLMs) or large multimodal models (LMMs). By leveraging the commonsense capabilities of LLMs and LMMs, these approaches (e.g., Holodeck[73], LEGENT[343]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models on high-volume, rich datasets has demonstrated remarkable capabilities and significant success in effectively addressing downstream applications. For instance, large language models such as ChatGPT, GPT-4, and LLaMA have not only excelled in the field of NLP but have also provided excellent problem-solving capabilities for downstream tasks. Therefore, is it possible to train an embodied large model in the robotics field, one that possesses strong generalization capabilities through training and can adapt to new scenarios and robotic tasks? This requires a large volume of embodied datasets to provide data for model training. Open X-Embodiment [285]: A collaboration of 21 institutions collected an embodied dataset from 22 different robots, showcasing 527 skills and 160,266 tasks. The data they collected consisted of authentic demonstration data from robots, obtained by recording the process of executing operations. This primarily focused on domestic and kitchen settings, involving items such as furniture, food, and tableware. The operations mainly centered around pick-and-place tasks, with a small portion involving more complex maneuvers. The high-capacity model RT-X trained on this dataset demonstrated excellent transfer capabilities. UMI [348]: proposed a data collection and policy learning framework. They designed a handheld gripper and an elegant interface for data collection, enabling portable, low-cost, and information-rich data collection for challenging bimanual and dynamic demonstration data. By simply modifying the training data, the robot can achieve zero-shot generalizable, bimanual, and precise tasks. Mobile ALOHA [349]: proposed Mobile ALOHA, a low-cost full-body mobile manipulation system. It can be used to collect task data for bimanual operations under full-body mobility, such as frying shrimp and serving dishes. Training agents with data collected by this system and static ALOHA can improve the performance of mobile manipulation tasks. Such agents can serve as home assistants or work assistants. In human-agent collaboration [350], humans and agents learn together during data collection, reducing human workload, accelerating data acquisition, and improving data quality. Specifically, in an embodied scenario, during data collection, humans provide initial action inputs. Subsequently, the agent refines these actions through iterative perturbation and denoising processes, gradually optimizing them to produce precise, high-quality operational demonstrations. This entire process can be summarized as follows: humans contribute intuition and diversity in operations, while agents handle optimization and stability, reducing reliance on operators, enabling execution of more complex tasks, and gathering higher-quality data.
@@ VII-B2 Simulated Data
The aforementioned data collection methods involve directly collecting demonstration data in the real world for agent training. Such collection methods often require significant manpower, material resources, and time, leading to inefficiency. Therefore, in most cases, researchers can choose to collect datasets in simulation environments for model training. Collecting data in simulation environments does not require extensive resources and can generally be automated by programs, saving a lot of time. CLIPORT [276], Transporter Networks [351]: They collected demonstration data from the Pybullet simulator for end-to-end network model training and successfully transferred the models from simulation to real world. GAPartNet [352]: constructed a large-scale part-centric interactive dataset GAPartNet, providing rich part-level annotations for perception and interaction tasks. They proposed a pipeline for domain-generalized 3D part segmentation and pose estimation, which can generalize well to unseen object categories in both simulators and the real world. SemGrasp [275]: built a large-scale grasping text-aligned dataset CapGrasp, a semantically rich dexterous hand grasping dataset from virtual environments.
VII-B3 Sim2Real Paradigms
@@
Figure 16: Five pipelines to achieve sim2real gap. “Sim2Real2Sim” reduces the gap by reconstructing real scenes. “TRANSIC” compensates for the sim2real transfer gap through human-corrected interventions. “Domain Randomization” enhances model transfer adaptability by simulating environmental diversity. “System Identification” improves sim-to-real environment similarity, thereby mitigating the sim-real gap. “Lang4Sim2Real” uses natural language to bridge two domains, learning invariant image representations and reducing visual gaps.
@@
Recently, several sim2real paradigms have been introduced. These methods aim to mitigate the need for extensive and costly real-world demonstration data by conducting extensive learning in simulation environments firstly, followed by migration to real-world settings. This section outlines five paradigms for sim2real transfer, as shown in Fig. 16.
Real2Sim2Real [353] is an approach that enhances imitation learning in real-world scenarios by leveraging reinforcement learning trained in a “digital twin” simulation environment. The method involves strengthening strategies through extensive RL within the simulation, followed by transferring these strategies to the real world to address data scarcity and achieve effective robotic operational imitation learning. Initially, existing methods such as Nerf and VR are used for scene scanning and reconstruction, and the constructed scene assets are imported into the simulator to achieve real-to-simulation fidelity. Subsequently, RL in the simulation fine-tunes the initial strategies derived from sparse expert demonstrations collected in the real world. Finally, the refined strategies are transferred to real-world settings. TRANSIC [354] is a novel idea that narrows the sim-to-real gap by enabling real-time human intervention to correct robot behaviors in real-world scenes. The method enhances sim-to-real transfer performance through a series of steps: initially, robots are trained using RL to establish foundational strategies within a simulation environment. Subsequently, these strategies are implemented on real robots, with humans intervening and correcting behaviors in real-time via remote control when errors occur. The data collected from these interventions are used to train a residual policy. Finally, integrating both foundational and residual policies ensures smoother trajectories in real-world applications following sim-to-real transfer. This approach significantly reduces the need for real-world data collection, thereby mitigating the burden while achieving sim2real. Domain Randomization [355, 356, 357] enhances the generalization of models trained in simulated environments to real-world scenarios by introducing parameter randomization during simulation. While both simulated and real environments involve perception through camera-acquired visual images, differences such as object friction and gloss make accurate simulation challenging. Therefore, by randomizing parameters during simulation training, a wide range of conditions can be covered, potentially encompassing variations that might occur in real-world settings. This approach boosts the robustness of trained models, enabling direct deployment from simulation to real environments. System Identification [358, 359] constructs an accurate mathematical model of physical scenes in real-world environments, encompassing parameters such as dynamics and visual rendering. This effort aims to make simulation environments closely resemble real-world settings, facilitating smooth transitions of models and strategies trained in simulation to real environments. Lang4Sim2Real [360] uses natural language as a bridge to address the gap between simulation and real-world environments, specifically by using textual descriptions of images as a cross-domain unified signal. This approach aids in learning domain-invariant image representations, thereby improving generalization performance across simulation and real environments. Initially, an encoder is pretrained on image data annotated with cross-domain language descriptions. Subsequently, using the learned domain-invariant representations, a multi-domain, multi-task language-conditioned behavioral cloning policy is trained. This method compensates for the scarcity of real-world data by leveraging additional information from abundant and inexpensive simulated data, thereby enhancing sim-to-real transfer.

### v3

@@ V-A3 Metrics
The model’s performance is usually assessed based on two aspects: navigation and question answering. In the navigation, many works adhere to the approach introduced by Das et al[265]. and utilize indicators like the distance to the target object upon completion of navigation ($d_{T}$), the change in distance to target from initial to final position ($d_{\varDelta}$) and the smallest distance to the target at any point in the episode ($d_{min}$) to evaluate the performance of the model. They are tested at 10, 30, or 50 actions away from the target. There are also works that measure it based on indicators such as trajectory length, intersection-over-union score for target object ($IoU$), etc. For question answering, the evaluation mainly involves mean rank ($MR$) of the ground-truth answer in the answer list and accuracy of the answers. Recently, Majumdar et al.[272] introduce the concept of an aggregate LLM-based correctness metric (LLM-Match) to evaluate the accuracy of open-vocabulary answers. Additionally, they assess efficiency by incorporating the normalized length of the agent’s path as a weight for the correctness metric.
@@ V-A4 Limitations
a) Dataset: Constructing datasets requires substantial manpower and resources. Additionally, there are still few large-scale datasets, and the metrics for evaluating model performance vary across different datasets, complicating the testing and comparison of performance, b) Model: Despite the advancements brought by LLMs, the performance of these models still lags significantly behind human levels. Future work may focus more on effectively storing environmental information explored by agents and guiding them to plan actions based on environmental memory and questions, while also enhancing the interpretability of their actions.
@@ V-B Embodied Grasping
TABLE X: Embodied grasping datasets.
@@
V-B1 Gripper
V-B2 Datasets
Recently, a substantial number of grasping datasets [287, 288, 289, 290, 291] have been generated. These datasets typically contain annotated grasping data based on images (RGB, depth), point clouds, or 3D scenes. With the advent of MLMs and the application of foundational language models to robotic grasping, there is an urgent need for datasets that include linguistic text. Consequently, existing datasets have been extended or reconstructed to create semantic-grasping datasets [293, 294, 299, 295]. These datasets are instrumental in studying grasping models grounded in language, enabling agents to develop a broad understanding of semantics.
@@ V-B3 Language-guided grasping
Spatial reasoning [293] refers to instructions that may include the spatial relationship of the object or part to be grasped, necessitating the inference of grasping posture based on the spatial relationships of objects within the scene. For example, “Grasp the keyboard that is to the right of the brown kleenex box” involves understanding and inferring the spatial arrangement of objects. Logical reasoning [294], on the other hand, involves instructions that may contain logical relationships requiring inference to discern human intent and subsequently grasp the target. For instance, “I am thirsty, can you give me something to drink?” would prompt the agent to potentially hand over a glass of water or a bottle of a beverage. The agent must ensure that the liquid does not spill during the handover, thus generating a reasonable grasping posture.
@@ V-B5 Modular Approaches
Figure 13: Chronological overview of the embodied agents. Different colors show different paradigms of embodied agents. MLM means the Multimodal Language Model which directly perceive the world and control the embodiment, VLM means the Visual-Language Model with the outer policy models, LLM + VLM means the LLM-based agent that perceives the world utilizing the VLM, and LLM means the Large-Language Model with visual context and outer policy models.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Initially, Symbolic Agents, rooted in symbolic reasoning, and Reactive Agents, known for their rapid responsiveness, were widely utilized. However, these agents were limited in handling complex strategies. Learning-based agents were subsequently developed to mitigate this limitation, yet they remained inadequate for large-scale real-world problems. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their capabilities from virtual space to the physical world, thereby becoming embodied agents [304].
@@
To enable embodied agents to operate in the information-rich and complex real world, the embodied agents have been developed to show strong multimodal perception, interaction and planning capabilities, as shown in Fig. 13. To complete a task, embodied agents typically involves the following process: 1) decomposing the abstract and complex task into specific subtasks, which is referred to as high-level Embodied Task Planning. 2) gradually implementing these subtasks by effectively utilizing Embodied Perception and Embodied Interaction models or leveraging the Foundation Model’s policy function, named low-level Embodied Action Planning. It is worth noting that task planning involves thinking before acting, and is therefore typically considered in cyber space. In contrast, action planning must account for effective interaction with the environment and feedback on this information to the task planner to adjust task planning. Thus, it is crucial for embodied agents to align and generalize their abilities from the cyber space to the physical world.
@@ VI-B1 Planning utilizing the Emergent Capabilities of LLMs
Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [325, 326, 327] . Interestingly, improving chain-of-thought (CoT) reasoning [328] is also a viable approach. ReAct [329] incorporates CoT reasoning with plan generation, providing more logical action generation. And some works utilize code as the reasoning medium instead of natural language. The Chain of Code method is particularly suitable for embodied agents in reasoning and executing tasks, where task planning is generated as code based on the available API library[330, 331, 332]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning, a focus of many LLM-based agent studies. For instance, Socratic Models [333] and Socratic Planner [334] use Socratic questioning to derive reliable planning.
@@ VI-B3 Planning utilizing the VLMs
In the EmbodiedGPT framework, the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning [310]. Unlike EmbodiedGPT, which directly uses third-person perspective images, LEO [347] encodes 2D egocentric images and 3D scenes into visual tokens. This method effectively perceives 3D world information and executes tasks accordingly. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [348]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [13, 308] and PaLM-E [307] to achieve alignment of visual and textual features in embodied scenarios. Additionally, innovative research such as Matcha [349] explores the integration of large language models (LLMs) with multimodal perception modules (e.g., weight, touch, sound) to enhance robot planning capabilities. For example, when executing the task “pick up a plastic block,” the robot can assess its weight by lifting it, determine if the sound it makes when tapped matches expected plastic sounds, or use tactile feedback to sense the block’s hardness or softness. Moreover, the VLP [350] model enhances planning by considering historical information and predicting future visual states. Once an LLM generates an action, the video model can simulate multiple potential video representations for each action. These video models then serve as heuristic functions to evaluate the proximity of each action to the goal, thereby improving planning effectiveness.
@@ VI-C2 Action utilizing VLA model
Different form previous approach that task planning and action execution are performed within the same system, this paradigm leverages the capabilities of embodied multimodal foundation models for planning and executing actions, reducing communication latency and improving system response speed and efficiency. In VLA models, the tight integration of perception, decision-making, and execution modules allows the system to handle complex tasks and adapt to changes in dynamic environments more efficiently. This integration also facilitates real-time feedback, enabling the agent to self-adjust strategies, thereby enhancing the robustness and adaptability of task execution [3, 310, 309]. However, this paradigm is undoubtedly more complex and costly, particularly when dealing with intricate or long-term tasks. Additionally, a key issue is that an action planner, without an embodied world model, cannot simulate physical laws using only the internal knowledge of an LLM. This limitation hinders the agent to accurately and effectively complete various tasks in the physical world, preventing the seamless transfer from cyber space to physical world.
@@ VII-A Embodied World Model
VII-A1 Generation-based Methods
TABLE XI: Summary of the embodied world methods discussed in VII-A.
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim-to-real approach[370], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen[364], Liu et al.[365], Lionel et al. [363]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with large language models (LLMs) or large multimodal models (LMMs). By leveraging the commonsense capabilities of LLMs and LMMs, these approaches (e.g., Holodeck[73], LEGENT[366]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models on high-volume, rich datasets has demonstrated remarkable capabilities and significant success in effectively addressing downstream applications. For instance, large language models such as ChatGPT, GPT-4, and LLaMA have not only excelled in the field of NLP but have also provided excellent problem-solving capabilities for downstream tasks. Therefore, is it possible to train an embodied large model in the robotics field, one that possesses strong generalization capabilities through training and can adapt to new scenarios and robotic tasks? This requires a large volume of embodied datasets to provide data for model training. Open X-Embodiment [309] is an embodied dataset from 22 different robots, showcasing 527 skills and 160,266 tasks. The data they collected consisted of authentic demonstration data from robots, obtained by recording the process of executing operations. This primarily focused on domestic and kitchen settings, involving items such as furniture, food, and tableware. The operations mainly centered around pick-and-place tasks, with a small portion involving more complex maneuvers. The high-capacity model RT-X trained on this dataset demonstrated excellent transfer capabilities. UMI [371] proposed a data collection and policy learning framework. They designed a handheld gripper and an elegant interface for data collection, enabling portable, low-cost, and information-rich data collection for challenging bimanual and dynamic demonstration data. By simply modifying the training data, the robot can achieve zero-shot generalizable, bimanual, and precise tasks. Mobile ALOHA [372] is a low-cost full-body mobile manipulation system. It can be used to collect task data for bimanual operations under full-body mobility, such as frying shrimp and serving dishes. Training agents with data collected by this system and static ALOHA can improve the performance of mobile manipulation tasks. Such agents can serve as home assistants or work assistants. In human-agent collaboration [373], humans and agents learn together during data collection, reducing human workload, accelerating data acquisition, and improving data quality. Specifically, in an embodied scenario, during data collection, humans provide initial action inputs. Subsequently, the agent refines these actions through iterative perturbation and denoising processes, gradually optimizing them to produce precise, high-quality operational demonstrations. This entire process can be summarized as follows: humans contribute intuition and diversity in operations, while agents handle optimization and stability, reducing reliance on operators, enabling execution of more complex tasks, and gathering higher-quality data.
@@ VII-B2 Simulated Data
The aforementioned data collection methods involve directly collecting demonstration data in the real world for agent training. Such collection methods often require significant manpower, material resources, and time, leading to inefficiency. Therefore, in most cases, researchers can choose to collect datasets in simulation environments for model training. Collecting data in simulation environments does not require extensive resources and can generally be automated by programs, saving a lot of time. CLIPORT [300] and Transporter Networks [374] collected demonstration data from the Pybullet simulator for end-to-end network model training and successfully transferred the models from simulation to real world. GAPartNet [375] constructed a large-scale part-centric interactive dataset GAPartNet, providing rich part-level annotations for perception and interaction tasks. They proposed a pipeline for domain-generalized 3D part segmentation and pose estimation, which can generalize well to unseen object categories in both simulators and the real world. SemGrasp [295] built a large-scale grasping text-aligned dataset CapGrasp, which is a semantically rich dexterous hand grasping dataset from virtual environments.
VII-B3 Sim-to-Real Paradigms
@@
Figure 17: Five pipelines to achieve sim-to-real gap. “Real2Sim2Real” reduces the gap by reconstructing real scenes. “TRANSIC” compensates for the sim-to-real transfer gap through human-corrected interventions. “Domain Randomization” enhances model transfer adaptability by simulating environmental diversity. “System Identification” improves sim-to-real environment similarity, thereby mitigating the sim-real gap. “Lang4Sim2Real” uses natural language to bridge two domains, learning invariant image representations and reducing visual gaps.
@@
Recently, several sim-to-real paradigms have been introduced, to mitigate the need for extensive and costly real-world demonstration data by conducting extensive learning in simulation environments, followed by migration to real-world settings. This section outlines five paradigms for sim-to-real transfer, as shown in Fig. 17.
Real2Sim2real [376] enhances imitation learning in real-world scenarios by leveraging reinforcement learning (RL) trained in a “digital twin” simulation environment. This method involves strengthening strategies through extensive RL within the simulation, followed by transferring these strategies to the real world to address data scarcity and achieve effective robotic operational imitation learning. Initially, NeRF and VR are used for scene scanning and reconstruction, and the constructed scene assets are imported into the simulator to achieve real-to-simulation fidelity. Subsequently, RL in the simulation fine-tunes the initial strategies derived from sparse expert demonstrations collected in the real world. Finally, the refined strategies are transferred to real-world settings.
TRANSIC [377] narrows the sim-to-real gap by enabling real-time human intervention to correct robot behaviors in real-world scenes. It enhances sim-to-real transfer performance through several steps: first, robots are trained using RL to establish foundational strategies within a simulation environment. Then, these strategies are implemented on real robots, with humans intervening and correcting behaviors in real-time via remote control when errors occur. The data collected from these interventions are used to train a residual policy. Integrating both foundational and residual policies ensures smoother trajectories in real-world applications following sim-to-real transfer. This approach significantly reduces the need for real-world data collection, thereby mitigating the burden while achieving effective sim-to-real transfer.
Domain Randomization [378, 379, 380] enhances the generalization of models trained in simulated environments to real-world scenarios by introducing parameter randomization during simulation. While both simulated and real environments involve perception through camera-acquired visual images, differences such as object friction and gloss make accurate simulation challenging. Randomizing parameters during simulation training covers a wide range of conditions, potentially encompassing variations that might occur in real-world settings. This approach boosts the robustness of trained models, enabling deployment from simulation to real environments.
System Identification [381, 382] constructs an accurate mathematical model of physical scenes in real-world environments, encompassing parameters such as dynamics and visual rendering. It aims to make simulation environments closely resemble real-world settings, facilitating smooth transitions of models trained in simulation to real environments.
Lang4sim-to-real [383] uses natural language as a bridge to address the sim-to-real gap by using textual descriptions of images as a cross-domain unified signal. This approach aids in learning domain-invariant image representations, thereby improving generalization performance across simulation and real environments. Initially, an encoder is pretrained on image data annotated with cross-domain language descriptions. Subsequently, using the learned domain-invariant representations, a multi-domain, multi-task language-conditioned behavioral cloning policy is trained. This method compensates for the scarcity of real-world data by leveraging additional information from abundant and inexpensive simulated data, thereby enhancing sim-to-real transfer.